1. If the radius of the circle increases at the rate of 0.01 cm per second, find the rate of change of the area when the radius is 3 cm long










1
Expert's answer
2021-10-13T06:51:58-0400

Solution.

We know that the area of the circle determine as "S=\u03c0r^2." Also we have "\frac{dr}{dt}=0.01" cm/sec, because the radius of the circle increases at the rate of 0.01 cm per second.

Find "\frac{dS}{dt}=2πr\frac{dr}{dt}." Such as r=3 cm, then "\frac{dS}{dt}=2π•3•0.01=0.06π" cm2/sec or 0.1884 cm2/sec.

Answer. 0.1884 cm2/sec